What is the frequency of a wave that has a speed of 0.4 m/s and a wavelength of 0.020 meters?
1
Expert's answer
2018-01-08T04:37:44-0500
solution

frequency=(wave speed)/(wavelenght)=0.4/0.02=20 Hz

answer 20Hz